OpenStudy (anonymous):

The perimeter of rectangle is 42 cm. The length is 7cm more than the width. find the length.

OpenStudy (anonymous):

P=2(l+w)=42 (l & w are length and width ) l-7=w can you solve it for l and w?

OpenStudy (anonymous):

my instructor got the answer 14 cm, but I dont know how

OpenStudy (anonymous):

42=2(l+l-7) 4l-14=42 4*l=42+14=56 l=56/4 l=14 (cm) here you go :)